# The Role of Graph Topology in the Performance of Biomedical Knowledge Graph Completion Models

### v1

@@ Appendix C Additional Results
Figure C.1: Mean reciprocal rank plotted against the average occurrence of edge cardinalities (a) and edge topological patterns (b) in the six datasets.
@@
Figure C.5: Distribution of Spearman-rank correlation between in-degree of same relation and how frequently the entity is incorrectly selected among the top-$100$ tail predictions, grouping test queries by relation type. A positive correlation means that KGE models are biased towards predicting entities with a larger number of incoming edges of the relation type considered in the query.
@@
(a) Median head out-degree and tail in-degree of same relation and number of unique relation tail entities.
@@
Figure C.8: Statistics of topological properties for the interaction types investigated in Section 5.3.
@@
Figure C.9: Demixing for the interaction types investigated in Section 5.3. For each test query, we compute how many of the top-$100$ predictions made by the model are contained in the set of entities used as tails by triples of the considered relations.
@@
Head out-degree s.r. $34$ $34$ $14$ $14$ $20$ $20$ $7937$ $7935$
Unique out-relations $4$ $6$ $4$ $20$ $4$ $8$ $3$ $3$
@@
Tail in-degree s.r. $164$ $193$ $208$ $208$ $27$ $27$ $20$ $15$
Unique in-relations $1$ $3$ $8$ $12$ $7$ $11$ $8$ $11$

### v2

@@ Appendix C Additional Results
Figure C.1: Hits@1 and Hits@10 on the test split achieved by the KGE models, for the six datasets.
Figure C.2: Effect of edge cardinality on Hits@10.
Figure C.3: Mean reciprocal rank (on the test set) plotted as a function of (a) the frequency of edge cardinalities (i.e., the fraction of triples in a dataset of a given edge cardinality, see Figure 4) and (b) the frequency of edge topological patterns (see Table 1).
@@
Figure C.7: Effect of head out-degree and tail in-degree on MRR, for the ConvE KGE model.
Figure C.8: Effect of the head and tail degrees of same relation type on Hits@10, for Hetionet and OpenBioLink.
Figure C.9: Distribution of Spearman-rank correlation between in-degree of same relation type and how frequently the entity is incorrectly selected among the top-$100$ tail predictions, grouping test queries by relation type. A positive correlation means that KGE models are biased towards predicting entities with a larger number of incoming edges of the relation type considered in the query.
@@
Figure C.11: The effect of having compositions on MRR, for the ConvE KGE model.
Figure C.12: The effect of having compositions on Hits@10, for Hetionet and OpenBioLink.
@@
(a) Median head out-degree and tail in-degree of same relation type and number of unique relation tail entities.
@@
Figure C.14: Statistics of topological properties for different interaction types.
@@
Figure C.15: Demixing for different interaction types. For each test query, we compute how many of the top-$100$ predictions made by the model are contained in the set of entities used as tails by triples of the considered relations.
Figure C.16: Hits@10 of the interaction types considered in Section 5.3.
@@
Head out-degree s.r. $34$ $34$ $14$ $14$ 2020 2020 $7937$ $7935$
Unique out-relations 44 66 44 2020 44 88 33 33
@@
Tail in-degree s.r. $164$ $193$ $208$ $208$ $27$ $27$ 2020 $15$
Unique in-relations 11 33 88 $12$ 77 $11$ 88 $11$